# State space representations of the Roesser type for convolutional layers

### v1

@@ 5 More general convolutions and examples
$\begin{bmatrix}x[i+1]\\ y[i]\end{bmatrix}=\begin{bmatrix}f&A&B\\ g&C&D\end{bmatrix}\begin{bmatrix}1\\ x[i]\\ u[i]\end{bmatrix}$ (52)
with state $x[i]\in\mathbb{R}^{n}$ and state dimension $n=rc_{\mathrm{in}}$ (Pauli et al., 2023a). Here, the matrices $A$, $B$, $C$, $D$, $f$, $g$ are
$f$ $=0,~{}$ $A=$ $\begin{bmatrix}0&I\\ 0&0\end{bmatrix},~{}$ $B$ $=\begin{bmatrix}0\\ I\end{bmatrix},$
@@
Remark 3
Theorem 2 requires $K[r_{1},r_{2}]$ to be invertible, which implies $c_{\mathrm{in}}=c_{\mathrm{out}}$ and that $K[r_{1},r_{2}]$ has full column rank. For $c_{\mathrm{in}}\neq c_{\mathrm{out}}$, the invertibility assumption is violated.
@@
$i_{1}$$i_{2}$$i_{1}$$i_{2}$
Figure 1: Visualization of state space representation for $3\times 3$ kernel with stride $s=1$ (left) and $s=2$ (right). At propagation step $i_{1},i_{2}$, the blue pixel input enters through $D$, the purple pixel input enters through $C_{2}$, the red pixel input enters through $B_{1}$, and the yellow pixel input enters through $A_{12}$.
@@
We consider a convolutional layer (1) with a $3\times 3$ kernel and stride $1$. A state space representation is given by (12) with
$\boldsymbol{A}=\left[\begin{array}[]{cc|cc}0&0&K[2,2]&K[2,1]\\ I&0&K[1,2]&K[1,1]\\ \hline\cr 0&0&0&I\\ 0&0&0&0\end{array}\right],~{}$ $\boldsymbol{B}$ $=\left[\begin{array}[]{c}K[2,0]\\ K[1,0]\\ \hline\cr 0\\ I\end{array}\right],$
$\boldsymbol{C}=\left[\begin{array}[]{cc|cc}0&I&K[0,2]&K[0,1]\end{array}\right],~{}$ $\boldsymbol{D}$ $=K[0,0],$
@@
We consider a convolutional layer (1) with a $2\times 3$ kernel and stride $1$. A state space representation (12) is given by
$\boldsymbol{A}=\left[\begin{array}[]{c|cc}0&K[1,2]&K[1,1]\\ \hline\cr 0&0&I\\ 0&0&0\end{array}\right],~{}$ $\boldsymbol{B}$ $=\left[\begin{array}[]{c}K[1,0]\\ \hline\cr 0\\ I\end{array}\right],$
$\boldsymbol{C}=\left[\begin{array}[]{c|cc}I&K[0,2]&K[0,1]\end{array}\right],~{}$ $\boldsymbol{D}$ $=K[0,0].$
@@
We consider a convolutional layer (1) with a $3\times 2$ kernel. A state space representation (12) is given by
$\boldsymbol{A}=\left[\begin{array}[]{cc|c}0&0&K[2,1]\\ I&0&K[1,1]\\ \hline\cr 0&0&0\end{array}\right],~{}$ $\boldsymbol{B}$ $=\left[\begin{array}[]{c}K[2,0]\\ K[1,0]\\ \hline\cr I\end{array}\right],$
$\boldsymbol{C}=\left[\begin{array}[]{cc|c}0&I&K[0,1]\end{array}\right],~{}$ $\boldsymbol{D}$ $=K[0,0].$
@@
Representing a convolution in state space requires the choice of a propagation direction for both dimensions. Usually, for image inputs we pick the upper left corner as the origin with $i_{1}=i_{2}=0$, cmp. Fig. 1. However, any other corner and corresponding propagation directions can also be chosen to represent the convolution equivalently. For recurrent layers (that in general are no FIR filters) the propagation dimension, i.e., time is predefined, and cannot be changed.
@@ 5.3 N-D convolutions
${\small\begin{bmatrix}K[r_{1},r_{2},r_{3}]&\cdots&K[r_{1},r_{2},0]&K[r_{1},r_{2}-1,r_{3}]&\cdots&K[r_{1},0,0]\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ K[0,r_{1},r_{3}]&\cdots&K[0,r_{2},0]&K[0,r_{2}-1,r_{3}]&\cdots&K[0,0,0]\end{bmatrix}}.$
@@
The mapping (14) can be represented as (28), where the matrices $f_{1},\ldots,D$ are given by
$\begin{split}&\left[\begin{array}[]{cccc}A_{12}&\cdots&A_{1d}&B_{1}\\ C_{2}&\cdots&C_{d}&D\end{array}\right]=\mathrm{mat}(K),\left[\begin{array}[]{ccc}A_{11}\\ C_{1}\end{array}\right]=\left[\begin{array}[]{ccc}0\\ I\end{array}\right],\\ &\left[\begin{array}[]{cccc}A_{22}&\dots&A_{2d}&B_{2}\\ \end{array}\right]=\left[\begin{array}[]{ccc}0&I\end{array}\right],\cdots,\left[\begin{array}[]{ccc}A_{dd}&B_{d}\\ \end{array}\right]=\left[\begin{array}[]{ccc}0&I\end{array}\right],\\ &A_{21}=0,\dots,A_{d1}=0,~{}A_{32}=0,\dots,A_{d2}=0,\dots,\\ &A_{dd-1}=0,~{}f_{1}=0,\dots,f_{d}=0,~{}g=b,\end{split}$
@@
The proof of Corollary 5.1 follows along the lines of the proof of Theorem 1 and is omitted. In the following, we provide an example.
@@
$\boldsymbol{A}\!=\!\left[\begin{array}[]{c|cc|c}0&K[1,1,1]&K[1,1,0]&K[1,0,1]\\ \hline\cr 0&0&0&I\\ 0&0&0&0\\ \hline\cr 0&0&0&0\end{array}\right]\!,$
$\boldsymbol{B}\!=\!\left[\begin{array}[]{c}K[1,0,0]\\ \hline\cr 0\\ I\\ \hline\cr I\end{array}\right]\!,~{}\boldsymbol{D}=K[0,0,0],$
$\boldsymbol{C}=\left[\begin{array}[]{c|cc|c}I&K[0,1,1]&K[0,1,0]&K[0,0,1]\end{array}\right],$
@@ 5.4 Dilated convolutions
Dilated convolutions insert spaces between the kernel elements. The number of skipped pixels/spaces is governed by an additional hyperparameter (Dumoulin and Visin, 2016). Expanding the filter allows to construct a state space representation for dilated convolutions. To illustrate the concept, we provide an example.
@@
5.5 Zero-padding
Zero-padding is achieved by choosing zero initial (and terminal) state conditions. Terminal state conditions are naturally zero for a signal in $\ell_{2e}^{c}(\mathbb{R}_{0}^{d})$. As zero-padding is common practice in neural networks, we choose zero initial conditions as the default in Definition 3.1 of the Roesser model. We distinguish between no zero-padding, same zero-padding, and full zero-padding, shown in Fig. 2. Full zero-padding includes padding of size $\boldsymbol{r}-1$ on all edges and same padding is chosen such that the input and the output dimensions remain the same. In case of no zero-padding and same zero-padding, the output signal needs to be cropped before entering the next layer, i.e., an image with support $[0,\boldsymbol{N}]$ is reduced in size by cropping the support to $[\boldsymbol{r},\boldsymbol{N}-\boldsymbol{r}]$ for no zero-padding or $[\mathrm{floor}(\boldsymbol{r}/2),\boldsymbol{N}-\mathrm{ceil}(\boldsymbol{r}/2)]$ for same zero-padding.
Figure 2: (a) No padding, (b) same padding, (c) full padding for a $3\times 3$ kernel (Dumoulin and Visin, 2016).
@@ 5.6 Strided convolutions
This means that we always shift the kernel by $s_{1},\ldots,s_{d}$ along the respective signal dimension $1,\ldots,d$.
Strided convolutions with stride $\boldsymbol{s}$ move the convolution kernel by $s_{1},\dots,s_{d}$ pixels in one propagation step over the input image (Dumoulin and Visin, 2016).
@@
Now we consider the case of a $3\times 3$ kernel and stride $s_{1}=s_{2}=2$ with input and output channels both $1$. To this end, we decompose the convolutional layer into an isometric isomorphism reshaping the input and a Roesser realization. The isomorphism is given by $\ell_{2e}^{1}(\mathbb{N}_{0}^{2})\to\ell_{2e}^{4}(\mathbb{N}_{0}^{2}),$
@@
$\boldsymbol{A}=\left[\begin{array}[]{c|cc}0&0&K[2,2]\\ \hline\cr 0&0&0\\ 0&0&0\end{array}\right],\qquad\boldsymbol{B}=\left[\begin{array}[]{cccc}0&0&K[2,1]&K[2,0]\\ \hline\cr 0&I&0&0\\ 0&0&0&I\end{array}\right],$
$\boldsymbol{C}=\left[\begin{array}[]{c|cc}I&K[1,2]&K[0,2]\end{array}\right],~{}$
@@
The convolution in Example 6 is visualized in Fig. 1 (right) at $i_{1},i_{2}$. For stride $s=2$, the reshaping operator, that we denote by $\mathrm{reshape}_{\boldsymbol{s}}$, vectorizes the $2\times 2$ input shown in blue into a $1\times 4$ vector. The reshaping operator is necessary, because a strided convolution is only shift invariant with respect to a shift by the stride $\boldsymbol{s}$ along $\boldsymbol{i}$ from $\mathcal{D}_{\mathrm{in}}$ to $\mathcal{D}_{\mathrm{out}}$. In case of $\boldsymbol{s}>1$, the shift-invariant input involves a batch of pixels and we use a reshape operation to vectorize this input. In its general form, $\mathrm{reshape}_{\boldsymbol{s}}$ is given by
@@
$\left[\begin{array}[]{c|c}\widetilde{A}_{12}&\widetilde{B}_{1}\\ \hline\cr\widetilde{C}_{2}&\widetilde{D}\end{array}\right]=\left[\begin{array}[]{ccc|ccc}K[r_{1},r_{2}]&\cdots&K[r_{1},s_{1}]&K[r_{1},s_{1}-1]&\cdots&K[r_{1},0]\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ K[s_{1},r_{2}]&\cdots&K[s_{1},s_{2}]&K[s_{1},s_{2}-1]&\cdots&K[s_{1},0]\\ \hline\cr K[s_{1}-1,r_{2}]&\cdots&K[s_{1}-1,s_{2}]&K[s_{1}-1,s_{2}-1]&\cdots&K[s_{1}-1,0]\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ K[0,r_{2}]&\cdots&K[0,s_{1}]&K[0,s_{2}-1]&\cdots&K[0,0]\\ \end{array}\right]$ (62)
In the following, we streamline how to obtain the matrices $\boldsymbol{A},\boldsymbol{B},\boldsymbol{C},\boldsymbol{D}$ for strided 2-D convolutions. For ease of exposition, we limit this part to 2-D convolutions, however, an extension to N-D convolutions is possible. To find the matrices $A_{12}$, $B_{1}$, $C_{2}$, and $D$, we reshape the blocks of (62) such that the dimension of $D$ matches the dimension of the reshaped vector input, i.e.,
@@
All remaining blocks are reshaped in the same fashion along the same dimensions and $\widetilde{A}_{12}$, $\widetilde{B}_{1}$ are padded with zeros if necessary (if $\operatorname{ceil}((r_{1}+1)/s_{1})\neq(r_{1}+1)/s_{1}$). The resulting state space representation for the reshaped signal has state, input and output dimensions $n_{1}=c_{\mathrm{out}}\operatorname{ceil}((r_{1}-s_{1}+1)/s_{1})$, $n_{2}=c_{\mathrm{in}}(r_{2}-s_{2}+1)s_{1}$, $n_{u}=s_{1}s_{2}c_{\mathrm{in}}$, $n_{y}=c_{\mathrm{out}}$. We note that in case $\boldsymbol{s}=\boldsymbol{r}$, $\boldsymbol{A}$, $\boldsymbol{B}$, $\boldsymbol{C}$ are empty matrices as the system has no memory. Finally, we present an additional example.
Example 7
We consider a convolutional layer (53) with a $5\times 5$ kernel and stride $2$. A possible state space representation (12) is given by
$\boldsymbol{A}\!=\!\left[\begin{array}[]{cc|cccccc}0&0&0&0&0&K[4,4]&K[4,3]&K[4,2]\\ I&0&K[3,4]&K[3,3]&K[3,2]&K[2,4]&K[2,3]&K[2,2]\\ \hline\cr 0&0&0&0&I&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&I\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\end{array}\right]\!,$
$\boldsymbol{B}=\left[\begin{array}[]{cccc}0&0&K[4,1]&K[4,0]\\ K[3,1]&K[3,0]&K[2,1]&K[2,0]\\ \hline\cr 0&0&0&0\\ I&0&0&0\\ 0&I&0&0\\ 0&0&0&0\\ 0&0&I&0\\ 0&0&0&I\end{array}\right]\!,~{}$
$\boldsymbol{C}=\left[\begin{array}[]{cc|cccccc}0&I&K[1,4]&K[1,3]&K[1,2]&K[0,4]&K[0,3]&K[0,2]\end{array}\right]\!,~{}$

### v2

@@ 5 General convolutions and examples
$\begin{bmatrix}x[i+1]\\ y[i]\end{bmatrix}=\begin{bmatrix}0&A&B\\ g&C&D\end{bmatrix}\begin{bmatrix}1\\ x[i]\\ u[i]\end{bmatrix}$ (24)
with state $x[i]\in\mathbb{R}^{n}$ and state dimension $n=rc_{\mathrm{in}}$ (Pauli et al., 2023a). Here, the matrices $A$, $B$, $C$, $D$, $g$ are
$A=$ $\begin{bmatrix}0&I_{(r-1)c_{\mathrm{in}}}\\ 0&0\end{bmatrix},~{}$ $B$ $=\begin{bmatrix}0\\ I_{c_{\mathrm{in}}}\end{bmatrix},$
@@
$i_{1}$$i_{2}$
$i_{1}$$i_{2}$
Figure 1: Visualization of state space representations for $3\times 3$ kernel with stride $\boldsymbol{s}=1$ (left) and $\boldsymbol{s}=2$ (right). At propagation step $i_{1},i_{2}$, the blue pixel input enters through $D$, the purple pixel input enters through $C_{2}$, the red pixel input enters through $B_{1}$, and the yellow pixel input enters through $A_{12}$.
@@
We consider a convolutional layer (1) with a $3\times 3$ kernel ($\boldsymbol{r}=2$) and stride $1$ ($\boldsymbol{s}=1$). A state space representation is given by (2) with
$\boldsymbol{A}=\left[\begin{array}[]{cc|cc}0&0&K[2,2]&K[2,1]\\ I_{c_{\mathrm{out}}}&0&K[1,2]&K[1,1]\\ \hline\cr 0&0&0&I_{c_{\mathrm{in}}}\\ 0&0&0&0\end{array}\right],~{}$ $\boldsymbol{B}$ $=\left[\begin{array}[]{c}K[2,0]\\ K[1,0]\\ \hline\cr 0\\ I_{c_{\mathrm{in}}}\end{array}\right],$
$\boldsymbol{C}=\left[\begin{array}[]{cc|cc}0&I_{c_{\mathrm{out}}}&K[0,2]&K[0,1]\end{array}\right],~{}$ $\boldsymbol{D}$ $=K[0,0],$
@@
We consider a convolutional layer (1) with a $2\times 3$ kernel ($r_{1}=1$, $r_{2}=2$) and stride $1$ ($\boldsymbol{s}=1$). A state space representation (2) is given by
$\boldsymbol{A}=\left[\begin{array}[]{c|cc}0&K[1,2]&K[1,1]\\ \hline\cr 0&0&I_{c_{\mathrm{in}}}\\ 0&0&0\end{array}\right],~{}$ $\boldsymbol{B}$ $=\left[\begin{array}[]{c}K[1,0]\\ \hline\cr 0\\ I_{c_{\mathrm{in}}}\end{array}\right],$
$\boldsymbol{C}=\left[\begin{array}[]{c|cc}I_{c_{\mathrm{out}}}&K[0,2]&K[0,1]\end{array}\right],~{}$ $\boldsymbol{D}$ $=K[0,0].$
@@
We consider a convolutional layer (1) with a $3\times 2$ kernel ($r_{1}=2$, $r_{2}=1$) and stride $1$ ($\boldsymbol{s}=1$). A state space representation (2) is given by
$\boldsymbol{A}=\left[\begin{array}[]{cc|c}0&0&K[2,1]\\ I_{c_{\mathrm{out}}}&0&K[1,1]\\ \hline\cr 0&0&0\end{array}\right],~{}$ $\boldsymbol{B}$ $=\left[\begin{array}[]{c}K[2,0]\\ K[1,0]\\ \hline\cr I_{c_{\mathrm{in}}}\end{array}\right],$
$\boldsymbol{C}=\left[\begin{array}[]{cc|c}0&I_{c_{\mathrm{out}}}&K[0,1]\end{array}\right],~{}$ $\boldsymbol{D}$ $=K[0,0].$
@@
Representing a convolution in state space requires the choice of a propagation direction for both dimensions. Usually, for image inputs we pick the upper left corner as the origin with $i_{1}=i_{2}=0$, cmp. Fig. 1. However, any other corner and corresponding propagation directions can also be chosen to represent the convolution equivalently. For general state space model layers the propagation dimension, i.e., time is predefined, and cannot be changed.
@@ 5.3 N-D convolutions
${\small\begin{bmatrix}K[r_{1},r_{2},r_{3}]&\cdots&K[r_{1},r_{2},0]&K[r_{1},r_{2}-1,r_{3}]&\cdots&K[r_{1},0,0]\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ K[0,r_{2},r_{3}]&\cdots&K[0,r_{2},0]&K[0,r_{2}-1,r_{3}]&\cdots&K[0,0,0]\end{bmatrix}}.$
@@
The mapping (19) can be represented as (20), where the matrices $A_{11},\ldots,D$ are given by
$\begin{split}&\left[\begin{array}[]{cccc}A_{12}&\cdots&A_{1d}&B_{1}\\ C_{2}&\cdots&C_{d}&D\end{array}\right]=\mathrm{mat}(K),\left[\begin{array}[]{ccc}A_{11}\\ C_{1}\end{array}\right]=\left[\begin{array}[]{ccc}0\\ I_{n_{1}}\end{array}\right],\\ &\left[\begin{array}[]{cccc}A_{22}&\dots&A_{2d}&B_{2}\\ \end{array}\right]=\left[\begin{array}[]{ccc}0&I_{n_{2}}\end{array}\right],\cdots,\left[\begin{array}[]{ccc}A_{dd}&B_{d}\\ \end{array}\right]=\left[\begin{array}[]{ccc}0&I_{n_{d}}\end{array}\right],\\ &A_{21}=0,\dots,A_{d1}=0,~{}A_{32}=0,\dots,A_{d2}=0,\dots,\\ &A_{dd-1}=0,~{}g=b,\end{split}$
@@
The proof of Corollary 5.1 follows along the lines of the proof of Theorem 1 and is omitted. Next, we provide an example.
@@
$\boldsymbol{A}\!=\!\left[\begin{array}[]{c|cc|c}0&K[1,1,1]&K[1,1,0]&K[1,0,1]\\ \hline\cr 0&0&0&I_{c_{\mathrm{in}}}\\ 0&0&0&0\\ \hline\cr 0&0&0&0\end{array}\right]\!,$
$\boldsymbol{B}\!=\!\left[\begin{array}[]{c}K[1,0,0]\\ \hline\cr 0\\ I_{c_{\mathrm{in}}}\\ \hline\cr I_{c_{\mathrm{in}}}\end{array}\right]\!,~{}\boldsymbol{D}=K[0,0,0],$
$\boldsymbol{C}=\left[\begin{array}[]{c|cc|c}I_{c_{\mathrm{out}}}&K[0,1,1]&K[0,1,0]&K[0,0,1]\end{array}\right],$
@@ 5.4 Dilated convolutions
Dilated convolutions insert spaces between the kernel elements. The number of skipped pixels/spaces is governed by an additional hyperparameter for dilation (Dumoulin and Visin, 2016). Expanding the filter allows to construct a state space representation for dilated convolutions. To illustrate the concept, we provide an example.
@@ 5.5 Strided convolutions
This means that, in each propagation step, we shift the kernel by $s_{1},\ldots,s_{d}$ steps along the respective signal dimension $1,\ldots,d$ (Dumoulin and Visin, 2016).
@@
Now we consider the case of a $3\times 3$ kernel ($r_{1}=r_{2}=2$) and stride $s_{1}=s_{2}=2$. To this end, we decompose the convolutional layer into an isometric isomorphism reshaping the input and a Roesser realization. The isomorphism is given by $\ell_{2e}^{1}(\mathbb{N}_{0}^{2})\to\ell_{2e}^{4}(\mathbb{N}_{0}^{2}),$
@@
$\boldsymbol{A}=\left[\begin{array}[]{c|cc}0&0&K[2,2]\\ \hline\cr 0&0&0\\ 0&0&0\end{array}\right],\qquad\boldsymbol{B}=\left[\begin{array}[]{cccc}0&0&K[2,1]&K[2,0]\\ \hline\cr 0&I_{c_{\mathrm{in}}}&0&0\\ 0&0&0&I_{c_{\mathrm{in}}}\end{array}\right],$
$\boldsymbol{C}=\left[\begin{array}[]{c|cc}I_{c_{\mathrm{out}}}&K[1,2]&K[0,2]\end{array}\right],~{}$
@@
The convolution in Example 6 is visualized in Fig. 1 (right) at $i_{1},i_{2}$. For stride $\boldsymbol{s}=2$, the reshaping operator, that we denote by $\mathrm{reshape}_{\boldsymbol{s}}$, vectorizes the $2\times 2$ input shown in blue into a $1\times 4$ vector. The reshaping operator is necessary, because a strided convolution is only shift invariant with respect to a shift by the stride $\boldsymbol{s}$ along $\boldsymbol{i}$ from $\mathcal{D}_{\mathrm{in}}$ to $\mathcal{D}_{\mathrm{out}}$. In case of $\boldsymbol{s}>1$, the shift-invariant input involves a batch of pixels and we use a reshape operation to vectorize this input. In its general form, $\mathrm{reshape}_{\boldsymbol{s}}$ is given by
@@
$\left[\begin{array}[]{c|c}\widetilde{A}_{12}&\widetilde{B}_{1}\\ \hline\cr\widetilde{C}_{2}&\widetilde{D}\end{array}\right]=\left[\begin{array}[]{ccc|ccc}K[r_{1},r_{2}]&\cdots&K[r_{1},s_{2}]&K[r_{1},s_{2}-1]&\cdots&K[r_{1},0]\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ K[s_{1},r_{2}]&\cdots&K[s_{1},s_{2}]&K[s_{1},s_{2}-1]&\cdots&K[s_{1},0]\\ \hline\cr K[s_{1}-1,r_{2}]&\cdots&K[s_{1}-1,s_{2}]&K[s_{1}-1,s_{2}-1]&\cdots&K[s_{1}-1,0]\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ K[0,r_{2}]&\cdots&K[0,s_{2}]&K[0,s_{2}-1]&\cdots&K[0,0]\\ \end{array}\right]$ (34)
@@
All remaining blocks are reshaped in the same fashion along the same dimensions and $\widetilde{A}_{12}$, $\widetilde{B}_{1}$ are padded with zeros if necessary (if $\operatorname{ceil}((r_{1}+1)/s_{1})\neq(r_{1}+1)/s_{1}$). The resulting state space representation for the reshaped signal has state, input and output dimensions $n_{1}=c_{\mathrm{out}}\operatorname{ceil}((r_{1}-s_{1}+1)/s_{1})$, $n_{2}=c_{\mathrm{in}}(r_{2}-s_{2}+1)s_{1}$, $n_{u}=s_{1}s_{2}c_{\mathrm{in}}$, $n_{y}=c_{\mathrm{out}}$. We note that in case $\boldsymbol{s}=\boldsymbol{r}$, $\boldsymbol{A}$, $\boldsymbol{B}$, $\boldsymbol{C}$ are empty matrices as the system has no memory.